Grade 12Algebra

If |z| = 3 , then the points representing the complex numbers -1+4z lie on : 1. Line 2. Circle 3. Parabola

Blue
10 Years agoGrade 12

1 Answer

jagdish singh singh
10 Years ago
\hspace{-0.60 cm}$Given $|z| = 3$ Now let $-1+4z=z`$(New Complex number)\\\\So $4z=z`+1\;,$ Then $4|z| = |z`+1|\;,$ So we get $|z`+1| = 12$\\\\Now put $z`=x`+iy`\;,$ We get $|x`+iy`+1| = 12$\\\\ So $(x`+1)^2+y`^2=12^2\;,$ Means it represent a Circle Whose \\\\ Center is at $(-1,0)$ and Radius $12$ unit.